The sum of two numbers is 43. The difference of the same two numbers is 12. What are the two numbers?
Answer by oberobic(2304)

Let x + y = 43.
So,
x = 43-y and
y = 43-x
.
x - y = 12
So,
(43-y) -y = 12
43 -2y = 12
-2y = -31
y = -31/-2 = 15.5
.
x + y = 43
x + 15.5 = 43
x = 43 - 15.5 = 27.5
.
Is the sum 43?
x = 27.5
y = 15.5
x + y = 43
Yes!
.
Is the difference 12?
x = 27.5
y = 15.5
x - y = 12
Yes!
.
So the answer is:
The two numbers are 15.5 and 27.5.
.
Done.